OpenStudy (anonymous):

what is the solution to 2 square root 15?

OpenStudy (anonymous):

2sqrt of 15 is essentially just the sqrt of 15, correct?

OpenStudy (lexilove13):

Is it like \[2\sqrt{15}\]

OpenStudy (anonymous):

7.74 is your answer if so, I can't show you any work because these types of problems require a calculator.

OpenStudy (anonymous):

But if my answer is wrong, could you please put your question in actual integer form? Like write the equation using the equation app right below the response box. :)

OpenStudy (lexilove13):

if the problem is set up like\[2\sqrt{15}\] I think you are trying to find what the entire square root is so the problem would be as so. \[2\sqrt{15} turns into \sqrt{4} * \sqrt{15} which equals \sqrt{60}\] then solve for the square of 60, which is 7.74

OpenStudy (anonymous):

thanks everyone for your help :3